Find the amount of energy required to raise the temperature of 9.19 g silicon from 18 oC to 58 oC

sSi = .71 J/goC


1
Expert's answer
2021-03-10T06:53:26-0500

Solution:

q = m × C × ΔT

q = m × C × (Tf - Ti),

where:

q = amount of heat energy gained or lost by substance

m = mass of sample

C = specific heat capacity (J oC-1 g-1)

Tf = final temperature

Ti = initial temperature

 

Thus:

q = (9.19 g) × (0.71 J oC-1 g-1) × (58 - 18)oC = 260.996 J = 261 J

q = 261 J


Answer: 261 Joules of energy required.